OpenStudy (anonymous):

solve each system of equations by the substitution method. Tell whether each system is inconsistent or consistent. If the system is consistent, tell whether it is independent or dependent 2x - y = 11 6x - 3y= -24

OpenStudy (bahrom7893):

substitution: \[2x-y=11\]\[-y=11-2x\]\[y=2x-11\]Plug that into the second equation

OpenStudy (anonymous):

how

OpenStudy (bahrom7893):

\[6x-3y=-24\]\[6x-3(2x-11)=-24\]Divide everything by 3:\[2x-(2x-11)=-8\]

OpenStudy (bahrom7893):

Simplify: \[-(-11)=-8\]\[11\ne-8\]

OpenStudy (bahrom7893):

Inconsistent

OpenStudy (anonymous):

can you help me with this one its hard for me. 2(x+y) -5(x+1)=3 -3x+2y=8

OpenStudy (anonymous):

Okay first expand this equation out and collect the like terms 2(x+y) -5(x+1)=3